OpenStudy (anonymous):

i dont understand this problem: 8^(x/2)=4^(x+1) anyone know?

OpenStudy (anonymous):

i ASSume you want to solve for x?

OpenStudy (anonymous):

i guess, the questions says to solve

OpenStudy (anonymous):

asks*

OpenStudy (anonymous):

right - this equation has x in the exponent (power) so we take logs of both sides: ln 8^(x/2) = ln 4^(x+1) by the laws of logarithms: (x/2) ln 8 = (x+1) ln 4 (x/2) = ((x+1) ln4) / ln 8) (x/2) / (x+1) = ln4) / ln 8) (x/2) / (x+1) = 2/3 excuse me - i must leave the room - be back later to complete it

OpenStudy (anonymous):

no problem. thanks for your reply in the meantime, jimmyrep

OpenStudy (anonymous):

so whats left is to solve this last equation 3x = 4(x+1) 3x = 4x + 4 x = -4

OpenStudy (anonymous):

the 3 main laws of logarithms are log x + log y = log xy log x - log y = log (x/y) log (x^y) = y log x

OpenStudy (anonymous):

thanks jimmyrep